Show that y=A sin (wt-kx) satisfies the equation d^2y/dc^2=1/v^2(d^2y/dt^2)


1
Expert's answer
2021-03-01T12:38:33-0500

"\frac{dy^2}{dx^2}=-Ak^2 \sin (wt-kx)\\
\frac{dy^2}{dt^2}=-Aw^2 \sin (wt-kx)\\v=\frac{w}{k}"

Thus,


"\frac{dy^2}{dx^2}=\frac{1}{v^2}
\frac{dy^2}{dt^2}"